A gas is separated from the air using an absorption column, which is randomly packed with 25mm metal Pall rings. Pure liquid solvent is used to reduce the mole fraction of gas A in air from 0.05 to 0.005 at a constant pressure and temperature. Total molar flow rates of entering gas and liquid to the column are 120 kmol/h and 180 kmol/h, respectively. The equilibrium relationship at operating conditions is !∗ = 1.2 ' and the gas side pressure drop in the column is selected as 200 Pa/m. By using the additional data given below calculate a) diameter of the column, b) percent of flooding gas velocity at operating conditions, c) height of the column, when height of the one overall liquid transfer unit is 0.95 m. MWG= 29 kg/kmol, MWL= 18 kg/kmol, ρG = 1.2 kg/m3 , ρL = 998 kg/m3 , µL= 1 cP
Answer #1

Given data

G = 120 kmol/h

L = 180 kmol/h

\rho _{g}= 1.2 Kg/m^{3}

\rho _{L}= 998 Kg/m^{3}

Viscosity = 1 cp

Metal pall rings of 25 mm is used

For metal pall rings from handbook

Packing factor (Fp) = 157 /m

The relationship between flooding velocity and pressure drop can be obtained in a graph from handbook

Pressure drop of column = 200 Pa/m

From handbook

816.404 Pa/m = 1 inch of water/ft

200 Pa/m = 0.2449 ~ 0.25 inch water /ft

X axis of graph

\frac{L}{G}\left ( \frac{\rho _{g}}{\rho _{L}} \right )^{0.5}

= 07月 180 12 120_998 = 0.0520

At X axis = 0.0520 and P = 0.25 inchesof water/ft

We get y axis from graph = 0.032

Y axis = \frac{G^{2}F_{p}\psi \mu ^{0.2}}{ \rho _{g}\rho _{L}g}

\psi = 1

0.032= \frac{G^{2}(157)(0.001)^{0.2}}{ (998)(1.2)(9.81)}

G = 3.0875 kg/m2s

G = (density of gas) (v)

Operating Velocity = (3.0875/1.2) = 2.5729 m/s

Molar flowrate of gas = 120 kmol/h

M. W of Gas = 29 Kg/kmol

Mass flowrate of gas(m) 29(120) = 3480 Kg/h

G = m/A

A = m/G = 3480/(3600×3.0875)

A = 0.31309 m2

A = π(d2) /4

0.31309 = π(d2) /4

d = 0.6313 m

Diameter of column = 0.6313 m

B)

From graph at X axis of 0.0520 , flooding velocity line has Y axis = 0.18

Substuting in Y axis equation we get

0.18= \frac{G^{2}(157)(0.001)^{0.2}}{ (998)(1.2)(9.81)}

G = 7.3227 kg/m2s

Flooding velocity (Vf) = G/(density of gas)

Vf = (7.3227/1.2) =6.10225 m/s

Operating velocity = 2.5729 m/s

V/Vf =( 2.5729/6.10225 ) = 0.42163

So operating velocity is 42.163% of flooding velocity

% of flooding = 42.163%

C)

HtoL= 0.95 m

Z = HtoL× NtoL

y1 = 0.05, Y1 = 0.05/(1+0.05) = 0.05263

y2 = 0.005 , Y2 = 0.005025

x2 = 0

L = Ls = 180 kmol/h

Gs = G(1-y1)

Gs = 120(1-0.05) = 114 kmol/h

Doing material balance

G_{s}(Y_{1}-Y_{2}) = L_{s}(X_{1}-X_{2})

114(0.05263-0.005025)= 180(X_{1}-0)

X1 = 0.030149

x1= 0.030149/(0.030149+1) = 0.029267

Equilibrium relationship

y = 1.2(x)

NtoL= (x2- x1) / (x-x*) lm

(x-x*) _{lm}= \frac{(x_{1}-x^{*}_{1})-(x_{2}-x_{2}*)}{ln\left ( \frac{x_{1}-x_{1}*}{x_{2}-x_{2}*} \right )}

y 1 is in equilibrium with x1

and y2 will be in equilibrium with x2

y = 1.2(x)

y1 = 0.05

y2 = 0.005

x 1 = 0.029267

x1* = 0.05/(1.2) = 0.04166

x2* = 0.005/(1.2) = 0.004166

(x-x*) _{lm}= \frac{(0.029267-0.04166)-(0-0.004166)}{ln\left ( \frac{0.029267-0.04166}{0-0.004166} \right )}

(x-x* ) lm= -7. 54649×10-3

NtoL = (x2 - x1) /(x-x*) lm

NtoL= (0-(0.029267) ) /(-7.54649×10-3)

NtoL = 3.8782

Z = HtoL × NtoL

Z = 0.95(3.8782) = 3.68429 m

Height of column = 3.68429 m

Please upvote if helpful